# Fast and accurate parameter estimation of high-redshift sources with the Einstein Telescope

### v1

@@ Abstract
The Einstein Telescope (ET), along with other third-generation gravitational wave (GW) detectors, will be a key instrument for detecting GWs in the coming decades. However, analyzing the data and estimating source parameters will be challenging, especially given the large number of expected detections – of order $10^{5}$ per year – which makes current methods based on stochastic sampling impractical. In this work, we use Dingo-IS to perform Neural Posterior Estimation (NPE) of high-redshift events detectable with ET in its triangular configuration. NPE is a likelihood-free inference technique that leverages normalizing flows to approximate posterior distributions. After training, inference is fast, requiring only a few minutes per source, and accurate, as corrected through importance sampling and validated against standard Bayesian inference methods. To confirm previous findings on the ability to estimate parameters for high-redshift sources with ET, we compare NPE results with predictions from the Fisher information matrix (FIM) approximation. We find that FIM underestimates sky localization errors substantially for most sources, as it does not capture the multimodalities in sky localization introduced by the geometry of the triangular detector. FIM also overestimates the uncertainty in luminosity distance by a factor of $\sim 3$ on average when the injected luminosity distance is $d^{\mathrm{inj}}_{\mathrm{L}}>10^{5}~{}$pc, further confirming that ET will be particularly well suited for studying the early Universe.
@@ I Introduction
The Einstein Telescope (ET) will be a leading instrument for detecting gravitational waves (GWs) in the next decades (Punturo et al., 2010; Maggiore et al., 2020; Branchesi et al., 2023; Abac et al., 2025), working alongside other third-generation (3G) GW observatories such as Cosmic Explorer (Reitze et al., 2019; Evans et al., 2023; Gupta et al., 2024) and LISA (Colpi et al., 2024).
ET will be able to detect GWs with unprecedented sensitivity, but this also brings great challenges to data analysis and parameter estimation (PE; Couvares et al., 2021). Some key issues include longer signals due to the wider frequency range, a large number of detections (of order $10^{5}$ events per year; Baibhav et al., 2019; Borhanian and Sathyaprakash, 2024; Maggiore et al., 2020; Branchesi et al., 2023), and overlapping signals (Samajdar et al., 2021). To address these challenges, new methods and technologies will be needed.
In the ET sensitivity band, a binary neutron star (BNS) signal can last from several minutes to hours. During this time, the noise characteristics of the detector may change (Zackay et al., 2021; Edy et al., 2021; Kumar et al., 2022), leading to non-stationary noise. Furthermore, transient noise signals, known as glitches, can overlap the signal in the data stream (Littenberg and Cornish, 2015; Cornish and Littenberg, 2015), as also occurs in current GW detectors. Adding to the challenge, with the advent of 3G detectors, manually removing glitches that coincide with signals will no longer be practical, highlighting the need for the development of new methods to systematically handle non-Gaussian noise (Legin et al., 2024; Raymond et al., 2024; Xiong et al., 2025).
Not only can glitches overlap with signals, but signals can also overlap with each other (Langendorff et al., 2023). This presents another potential issue for data analysis in the ET era (Wu and Nitz, 2023; Janquart et al., 2023). Parameter estimation becomes biased when the merger times of two overlapping signals are close, and accuracy worsens when their signal-to-noise ratios (SNRs) are similar (Samajdar et al., 2021; Pizzati et al., 2022; Relton and Raymond, 2021; Himemoto et al., 2021; Antonelli et al., 2021; Wang et al., 2024).
The parameters $\theta$ describing a GW signal are estimated from data $d$ using Bayes’ theorem (van der Sluys et al., 2008, 2009; Thrane and Talbot, 2019; Christensen and Meyer, 2022):
@@
where $\mathcal{L}(d|\theta)$ is the likelihood function, $\pi(\theta)$ is the prior distribution and $\mathcal{Z}(d)=\int\mathrm{d}\theta\mathcal{L}(d|\theta)\pi(\theta)$ the evidence. The posterior distribution $p(\theta|d)$ is typically represented by a set of samples obtained with stochastic sampling methods (Ashton et al., 2019).
@@
Several methods have been developed to accelerate stochastic sampling, including relative binning (also called heterodyning; Cornish, 2010; Veitch et al., 2015; Leslie et al., 2021), multibanding (or adaptive frequency resolution; Vinciguerra et al., 2017; Aubin et al., 2021; Morisaki, 2021; Alléné et al., 2025), and reduced-order quadrature (Canizares et al., 2015; Smith et al., 2021; Tissino et al., 2023). All of these aim to simplify likelihood evaluations and, therefore, accelerate convergence. These techniques have already been tested and shown to be effective for 3G detectors, though under certain simplifying assumptions (Smith et al., 2021), such as ignoring Earth’s rotation for long-duration signals (Narola et al., 2024; Hu and Veitch, 2024).
@@
Recent years have seen significant advances in machine learning-based analysis methods, resulting in substantial speedup in PE with precision comparable to traditional methods (Cuoco et al., 2021; Chua and Vallisneri, 2020; Gabbard et al., 2022; Zhao et al., 2023; Bhardwaj et al., 2023; Cuoco et al., 2024; Hu et al., 2024; Polanska et al., 2024; Stergioulas, 2024).
@@
One such method is simulation-based inference (or likelihood-free inference; Greenberg et al., 2019; Cranmer et al., 2020), which bypasses the need for likelihood computation by leveraging the ability to generate data through simulations. The core idea in the GW context is to approximate the posterior distribution $p(\theta|d)$ in Eq. 1 using normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2021). These are invertible transformations parametrized by neural networks trained to transform a multivariate Gaussian base distribution into the target posterior distribution. This approach is also known as neural posterior estimation (NPE; Papamakarios and Murray, 2016). Several previous studies have applied NPE or related methods in the context of GWs (Chatterjee et al., 2019; Shen et al., 2022; Delaunoy et al., 2020; Green et al., 2020; Green and Gair, 2021; Krastev et al., 2021; Dax et al., 2021, 2023; Wildberger et al., 2023; Bhardwaj et al., 2023; Kolmus et al., 2024; Mould et al., 2025).
@@
Normalizing flows can be used to accelerate stochastic sampling by improving the efficiency of sample proposals. For instance, they are integrated into nested sampling with nessai (Williams et al., 2021, 2023), significantly reducing the number of required likelihood evaluations (Speagle, 2020). Similarly, normalizing flows can enhance Markov Chain Monte Carlo (MCMC) methods, as implemented in jim (Wong et al., 2023a; Wouters et al., 2024), which relies on flowMC (Wong et al., 2023b) and ripple to speed up waveform generation with jax (Edwards et al., 2024; Bradbury et al., 2018).
@@
Additionally, NPE can accelerate population inference (Leyde et al., 2024). Extracting population properties from a large set of detected events $\{d\}$ using hierarchical Bayesian analysis can be more computationally demanding than estimating parameters for individual events. Resolving this issue with new methods is crucial and will be necessary before the beginning of the 3G detector era (Talbot et al., 2019; Talbot, 2021; Mastrogiovanni et al., 2024; Meesum Qazalbash, 2024; Colloms et al., 2025).
@@
The methods outlined above were developed primarily for current GW detectors. However, they hold significant promise for 3G detectors as well (Hu and Veitch, 2024).
@@
In this decade, it is crucial to shape the design of 3G detectors to maximize their scientific output (Branchesi et al., 2023). To guide decisions, large simulated populations of compact binary coalescences have been analyzed using the Fisher information matrix (FIM) approximation (Dupletsa et al., 2023; de Souza and Sturani, 2023; Iacovelli et al., 2022; Gupta et al., 2024). However, it is now necessary to develop new methods that can perform full PE for sources detectable by 3G detectors.
@@
This study has two main objectives. First, we apply and test NPE to provide full PE for sources detectable with the ET in its triangular configuration. We validate NPE by comparing its output with results obtained with standard inference libraries, such as Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020). Second, we validate previous results on ET performance in estimating source parameters by comparing results from NPE and FIM, by running a large-scale injection campaign. While previous studies have examined uncertainties in sky localization and other parameters using standard Bayesian inference and FIM for second-generation detectors, like LIGO, Virgo (Grover et al., 2014; Rodriguez et al., 2013; Dupletsa et al., 2025), and LISA (Porter and Cornish, 2015), we are not aware of any that have done this thoroughly for ET.
@@
We will focus on sources that merge at moderate and high redshift (i.e., $1\lesssim z\lesssim 45$) for several reasons. First, these signals are short in the ET frequency band, lasting less than a few seconds, making them more accessible for training NPE. Second, high-redshift signals have low SNRs, making reliable PE using the FIM challenging (Vallisneri, 2008). Third, only 3G GW detectors are sensitive to high-redshift sources, which makes these sources a unique component of their scientific case.
@@
Several physical processes lead to the formation of high-redshift sources in our Universe. For example, Population III (Pop. III) stars, which form from metal-free gas at $z>20$ (Haiman et al., 1996; Tegmark et al., 1997; Abel et al., 2002; Yoshida et al., 2006; Klessen and Glover, 2023), are believed to have been the first stars to form and are thought therefore to produce the earliest stellar-origin black holes (Kinugawa et al., 2016; Hartwig et al., 2016; Belczynski et al., 2017; Tanikawa et al., 2022a; Costa et al., 2023; Nandal et al., 2023; Liu et al., 2024a, b; Tanikawa, 2024; Mestichelli et al., 2024). As a result, Pop. III binary black holes (BBHs) have been extensively studied in recent years as sources of high-redshift GWs (Kinugawa et al., 2020; Liu and Bromm, 2020a, b; Tanikawa et al., 2022b, 2021; Wang et al., 2022). In particular, Santoliquido et al. (2023) found that over 20% of detectable Pop. III BBHs can occur at $z>8$. Additionally, primordial black holes are also expected to form at high redshift, covering a wide range of masses and merger rates (De Luca et al., 2020; De Luca, 2022; Franciolini et al., 2022; Ng et al., 2022a, b, 2021).
@@
This manuscript is organized as follows: Sec. II describes the methodology, including detector and waveform parametrization (Sec. II.1), chosen priors (Sec. II.2), NPE (Sec. II.4), nested sampling with Bilby (Sec. II.3), and the FIM approximation (Sec. II.5). Results for a single source and population analysis are presented in Secs. III.1 and III.2, respectively. The degeneracy in sky localization introduced by the triangular shape of ET is discussed in Sec. IV.1, followed by the impact of higher-order modes and precessing spins on our results (Sec. IV.2). In Sec. IV.3, we estimate the parameters of injected sources based on an astrophysical population of BBH mergers formed from Pop. III stars, while implications of our work on previous findings are discussed in Sec. IV.4. Sec. IV.5 compares the estimated energy required for NPE and stochastic sampling to analyze a population of sources. We discuss caveats of our methods in Sec. IV.6 and conclude in Sec. V.
@@ II.1 Detector and waveforms
We consider a single triangular ET detector, hereafter referred to as ET-$\Delta$, with an arm length of 10 km, located in Sardinia, which is a candidate site (Branchesi et al., 2023). Fig. 1 shows the locations of the three independent detectors that make up ET-$\Delta$, hereafter referred to as ET-1, ET-2, and ET-3, respectively. The triangular configuration is not the only proposed design for ET. Another option is the 2L configuration (Branchesi et al., 2023), which consists of two L-shaped detectors with 15 km arms located in Europe.
For our analysis, we use the HFLF cryogenic sensitivity curve (ET collaboration, 2023), which includes both a high-frequency instrument and a cryogenic low-frequency instrument. The estimated noise amplitude spectral density (ASD) is shown in Fig. 2. We also consider a frequency range between $f_{\mathrm{min}}=6$ Hz and $f_{\mathrm{max}}=256$ Hz. We use 8-second time segments, resulting in frequency bins of size $\mathrm{d}f=0.125$ Hz.
The merger frequency, estimated as the frequency at which the inspiral phase ends $f_{\mathrm{ISCO}}$, is given by Eq. 14.108 in Maggiore (2018),
@@
In a first approximation, the signal duration $\tau$ scales with the lower cutoff frequency $f_{\mathrm{min}}$ as given by Eq. 4.21 of Maggiore (2007):
@@
where $\mathcal{M}_{\mathrm{s}}$ is the source-frame chirp mass. By setting the frequency limit to $f_{\mathrm{min}}=6$ Hz, we ensure that the duration of all processed signals remains at most 8 seconds. However, this choice may slightly underestimate ET’s parameter estimation capabilities, whose sensitivity is expected to extend down to $\sim 2~{}\mathrm{Hz}$ (see Sec. IV.6 for further discussion).
We define spin-aligned quasi-circular waveforms with parameters $\theta$ = $\{\mathcal{M}_{\mathrm{d}}$, $q$, $d_{\mathrm{L}}$, $\mathrm{dec}$, $\mathrm{ra}$, $\theta_{\mathrm{jn}}$, $\psi$, $t_{\mathrm{geocent}}$, $\chi_{1}$, $\chi_{2}$, $\phi_{c}\}$, where $\mathcal{M}_{\mathrm{d}}$, $q$, and $d_{\mathrm{L}}$ denote the detector-frame chirp mass, mass ratio, and luminosity distance, respectively; ra and dec are the sky position coordinates, $\theta_{\mathrm{jn}}$ is the angle between the the line of sight and the total angular momentum $\vec{J}$, $\psi$ is the polarization angle, $t_{\mathrm{geocent}}$ is the coalescence time at the center of the Earth, and $\phi_{c}$ is the phase at coalescence. The waveforms also depend on the aligned spin parameters $\chi_{1}$ and $\chi_{2}$, defined as $\chi_{i}=a_{i}\cos\theta_{i}$, where $a_{i}=|\chi_{i}|$ is the spin magnitude and $\theta_{i}$ is the angle with respect to the orbit angular momentum, defined as:
$\theta_{i}=\begin{cases}0,&\text{if }\chi_{i}>0,\\ \pi,&\text{if }\chi_{i}<0\,.\\ \end{cases}$ (4)
We use frequency-domain waveforms, fixing the reference frequency to $f_{\mathrm{ref}}=20$ Hz (Varma et al., 2022), with the waveform approximant IMRPhenomXPHM (Pratten et al., 2021), which models subdominant modes $[(\ell,|m|)=(2,1),(3,3),(3,2),(4,4)]$ in addition to the dominant mode $[(\ell,|m|)=(2,2)]$. A detailed discussion on higher-order modes can be found in Sec. IV.2.

### v2

@@ Abstract
The Einstein Telescope (ET), along with other third-generation gravitational wave (GW) detectors, will be a key instrument for detecting GWs in the coming decades. However, analyzing the data and estimating source parameters will be challenging, especially given the large number of expected detections–on the order of $10^{5}$ per year–which makes current methods based on stochastic sampling impractical. In this work, we use Dingo-IS to perform neural posterior estimation (NPE) of high-redshift events detectable with ET in its triangular configuration. NPE is a likelihood-free inference technique that leverages normalizing flows to approximate posterior distributions. After training, inference is fast, requiring only a few minutes per source, and accurate, as corrected through importance sampling and validated against standard Bayesian inference methods. To confirm previous findings on the ability to estimate parameters for high-redshift sources with ET, we compare NPE results with predictions from the Fisher information matrix (FIM) approximation. We find that NPE correctly recovers the eight degenerate sky modes induced by the triangular detector geometry, missed by the FIM analysis, resulting in an underestimation of sky localization uncertainties for most sources. FIM also overestimates the uncertainty in luminosity distance by a factor of $\sim 3$ on average when the injected luminosity distance is $d^{\mathrm{inj}}_{\mathrm{L}}>10^{5}$~Mpc, further confirming that ET will be particularly well suited for studying the early Universe.
@@ I Introduction
The Einstein Telescope (ET) will be a leading instrument for detecting gravitational waves (GWs) in the next decades (Punturo:2010zz; ET:2019dnz; Branchesi:2023mws; Abac:2025saz), working alongside other third-generation (3G) GW observatories such as Cosmic Explorer (Reitze:2019iox; Evans:2023euw; Gupta:2023lga) and LISA (LISA:2024hlh).
ET will be able to detect GWs with unprecedented sensitivity, but this also brings great challenges to data analysis and parameter estimation (PE; Couvares:2021ajn). Some key issues include longer signals due to the wider frequency range, a large number of detections (of order $10^{5}$ events per year; Baibhav:2019gxm; 2024PhRvD.110h3040B; ET:2019dnz; Branchesi:2023mws), and overlapping signals (2021PhRvD.104d4003S). To address these challenges, new methods and technologies will be needed.
In the ET sensitivity band, a binary neutron star (BNS) signal can last from several minutes to hours. During this time, the noise characteristics of the detector may change (2021PhRvD.104f3034Z; 2021PhRvD.103l4061E; 2022arXiv220212762K), leading to non-stationary noise. Furthermore, transient noise signals, known as glitches, can overlap the signal in the data stream (2015PhRvD..91h4034L; 2015CQGra..32m5012C), as also occurs in current GW detectors. Adding to the challenge, with the advent of 3G detectors, manually removing glitches that coincide with signals will no longer be practical, highlighting the need for the development of new methods to systematically and automatically handle non-Gaussian noise (Narola:2024qdh; Legin:2024gid; Raymond:2024xzj; 2025PhRvD.111b4019X; Sun:2023vlq).
Not only can glitches overlap with signals, but signals can also overlap with each other (2023PhRvL.130q1402L). This presents another potential issue for data analysis in the ET era (Wu:2022pyg; Janquart:2022fzz). Parameter estimation becomes biased when the merger times of two overlapping signals are close, and accuracy worsens when their signal-to-noise ratios (SNRs) are similar (2021PhRvD.104d4003S; 2022PhRvD.105j4016P; 2021PhRvD.104h4039R; 2021PhRvD.104d4010H; 2021MNRAS.507.5069A; 2024CQGra..41e5011W; Papalini:2025exy).
The parameters $\theta$ describing a GW signal are estimated from data $d$ using Bayes’ theorem (vanderSluys:2007st; vanderSluys:2009bf; Thrane:2018qnx; Christensen:2022bxb):
@@
where $\mathcal{L}(d|\theta)$ is the likelihood function, $\pi(\theta)$ is the prior distribution and $\mathcal{Z}(d)=\int\mathrm{d}\theta\mathcal{L}(d|\theta)\pi(\theta)$ is the evidence. The posterior distribution $p(\theta|d)$ is typically represented by a set of samples obtained with stochastic sampling methods (Ashton:2018jfp).
@@
Several methods have been developed to accelerate stochastic sampling, including relative binning (also called heterodyning; 2010arXiv1007.4820C; Veitch:2014wba; 2021PhRvD.104l3030L), multibanding (or adaptive frequency resolution; Vinciguerra:2017ngf; Aubin:2020goo; 2021PhRvD.104d4062M; Allene:2025saz), and reduced-order quadrature (2015PhRvL.114g1104C; 2021PhRvL.127h1102S; 2023PhRvD.107h4037T). All of these aim to simplify likelihood evaluations and, thus accelerate convergence. These techniques have already been tested and shown to be effective for 3G detectors (Baral:2025geo), though under certain simplifying assumptions (Smith:2021bqc), such as ignoring Earth’s rotation for long-duration signals (2024PhRvD.110h4085N; 2024arXiv241202651H).
@@
Recent years have seen significant advances in machine-learning-based analysis methods, resulting in substantial speedup in PE with precision comparable to traditional methods (Cuoco:2020ogp; 2020PhRvL.124d1102C; 2022NatPh..18..112G; 2023arXiv231115585Z; Bhardwaj:2023xph; 2024arXiv241215046C; 2024arXiv241203454H; 2024arXiv241021076P; 2024arXiv240107406S).
@@
One such method is simulation-based inference (or likelihood-free inference; DBLP:journals/corr/abs-1905-07488; Cranmer_2020), which bypasses the need for likelihood computation by leveraging the ability to generate data through simulations. The core idea in the GW context is to approximate the posterior distribution $p(\theta|d)$ in Eq. 1 using normalizing flows (Rezende:2015ocs; Papamakarios:2019fms; Kobyzev_2021). These are invertible transformations parametrized by neural networks trained to transform a multivariate Gaussian base distribution into the target posterior distribution. This approach is also known as neural posterior estimation (NPE; Papamakarios:2016ctj). Several previous studies have applied NPE or related methods in the context of GWs (Chatterjee:2019gqr; Shen:2019vep; 2020arXiv201012931D; 2020PhRvD.102j4057G; Green:2020dnx; Krastev:2020skk; 2021PhRvL.127x1103D; 2023PhRvL.130q1403D; 2023mla..confE..34W; 2023PhRvD.108d2004B; 2024arXiv240302443K; Mould:2025dts; Papalini:2025exy).
@@
Normalizing flows can be used to accelerate stochastic sampling by improving the efficiency of sample proposals. For instance, they are integrated into nested sampling with nessai (2021PhRvD.103j3006W; 2023MLS&T...4c5011W), significantly reducing the number of required likelihood evaluations (2020MNRAS.493.3132S). Similarly, normalizing flows can enhance Markov Chain Monte Carlo (MCMC) methods, as implemented in jim (2023ApJ...958..129W; 2024PhRvD.110h3033W), which relies on flowMC (2023JOSS....8.5021W) and ripple to speed up waveform generation and enable automatic differentiation with jax (2024PhRvD.110f4028E; jax2018github).
@@
Additionally, NPE can accelerate population inference (2024PhRvD.109f4056L). Extracting population properties from a large set of detected events $\{d\}$ using hierarchical Bayesian analysis can be more computationally demanding than estimating parameters for individual events. Resolving this issue with new methods is crucial and will be necessary before the beginning of the 3G detector era (2019PhRvD.100d3030T; gwpop_pipe; Mastrogiovanni:2023zbw; gwkokab2024github; Colloms:2025hib; Mancarella:2025uat; Mould:2025dts).
@@
The methods outlined above were developed primarily for current GW detectors. However, they hold significant promise for 3G detectors as well (2024arXiv241202651H).
@@
In this decade, it is crucial to shape the design of 3G detectors to maximize their scientific output (Branchesi:2023mws). To guide decisions, large simulated populations of compact binary coalescences have been analyzed using the Fisher information matrix (FIM) approximation (Dupletsa:2022scg; 2023A&C....4500759D; 2022ApJS..263....2I; Gupta:2023lga; Borhanian:2022czq; Begnoni:2025oyd). However, it is now necessary to develop new methods that can perform full PE for sources detectable by 3G detectors.
@@
This study has two main objectives. First, we apply and test NPE to provide full PE for sources detectable with the ET in its triangular configuration. We validate NPE by comparing its output with results obtained with standard inference libraries, such as Bilby (Ashton:2018jfp; Romero-Shaw:2020owr). Second, we validate previous results on ET performance in estimating source parameters by comparing results from NPE and FIM, by running a large-scale injection campaign. While previous studies have examined uncertainties in sky localization and other parameters using standard Bayesian inference and FIM for second-generation detectors, like LIGO, Virgo (Grover:2013sha; Rodriguez:2013mla; Dupletsa:2024gfl), and LISA (Porter:2015eha), we are not aware of any that have done this thoroughly for ET.
@@
We will focus on sources that merge at moderate and high redshift (i.e., $1\lesssim z\lesssim 45$) for several reasons. First, these signals are short in the ET frequency band, lasting less than a few seconds, making them more accessible for training NPE. Second, high-redshift signals have low SNRs, making reliable PE using the FIM challenging (2008PhRvD..77d2001V). Third, detecting and characterizing high-redshift sources is central to the scientific goals of 3G GW detectors (Abac:2025saz), highlighting the need for methods that can provide fast and accurate PE.
@@
Several physical processes lead to the formation of high-redshift sources in our Universe. For example, Population III (Pop. III) stars, which form from metal-free gas at $z>20$ (Haiman1996; Tegmark1997; abel2002; yoshida2006; klessen2023), are believed to have been the first stars to form and are thought therefore to produce the earliest stellar-origin black holes (kinugawa2016; Hartwig2016; belczynski2017; tanikawa2022b; costa2023; Nandal2023; liu2023arXiv; Boyuan2024; tanikawa2024; Mestichelli:2024djn). As a result, Pop. III binary black holes (BBHs) have been extensively studied in recent years as sources of high-redshift GWs (kinugawa2020; liubromm2020GW190521; liubromm2020; tanikawa2022; tanikawa2021b; wang2022; Plunkett:2025mjr). In particular, santoliquido2023 found that over 20% of detectable Pop. III BBHs can occur at $z>8$. Additionally, primordial black holes are expected to form at high redshift, covering a wide range of masses and merger rates (DeLuca:2020qqa; DeLuca:2022iix; Franciolini:2022tfm; Ng:2021sqn; Ng:2022agi; Ng:2020qpk).
@@
This manuscript is organized as follows: Sec. II describes the methodology, including detector and waveform parametrization (Sec. II.1), chosen priors (Sec. II.2), NPE (Sec. II.4), nested sampling with Bilby (Sec. II.3), and the FIM approximation (Sec. II.5). Results for a single source and population analysis are presented in Secs. III.1 and III.2, respectively. The degeneracies in sky localization (Sec. IV.1) and geocentric time (Sec. IV.2) introduced by the triangular shape of ET are discussed first, followed by an analysis of the impact of higher-order modes and precessing spins on our results (Sec. IV.3). In Sec. IV.4, we estimate the parameters of injected sources based on an astrophysical population of BBH mergers formed from Pop. III stars, while implications of our work on previous findings are discussed in Sec. IV.5. Section IV.6 compares the estimated energy required for NPE and stochastic sampling to analyze a population of sources. We discuss caveats of our methods in Sec. IV.7 and conclude in Sec. V.
@@ II.1 Detector and waveforms
We consider a single triangular ET detector, hereafter referred to as ET-$\Delta$, with an arm length of 10 km, located in Sardinia, which is a candidate site (Branchesi:2023mws). Figure 1 shows the locations of the three independent detectors that make up ET-$\Delta$, hereafter referred to as ET-1, ET-2, and ET-3, respectively. The triangular configuration is not the only proposed design for ET. Another option is the 2L configuration (Branchesi:2023mws), which consists of two L-shaped detectors with 15 km arms located in Europe.
For our analysis, we use the HFLF cryogenic sensitivity curve (psd), which includes both a high-frequency instrument and a cryogenic low-frequency instrument. The estimated noise amplitude spectral density (ASD) is shown in Fig. 2. We also consider a frequency range between $f_{\mathrm{min}}=6$ Hz and $f_{\mathrm{max}}=256$ Hz. We use eight-second time segments, resulting in frequency bins of size $\mathrm{d}f=0.125$ Hz.
The merger frequency, estimated as the frequency at which the inspiral phase ends $f_{\mathrm{ISCO}}$, is given by Eq. 14.108 in 2018gwv..book.....M,
@@
In a first approximation, the signal duration $\tau$ scales with the lower cutoff frequency $f_{\mathrm{min}}$ as given by Eq. 4.21 of Maggiore:2007ulw:
@@
where $\mathcal{M}_{\mathrm{s}}$ is the source-frame chirp mass. By setting the frequency limit to $f_{\mathrm{min}}=6$ Hz, we ensure that the duration of all processed signals remains at most eight seconds. However, this choice may slightly underestimate ET’s parameter estimation capabilities, whose sensitivity is expected to extend down to $\sim 2~\mathrm{Hz}$ (see Sec. IV.7 for further discussion).
We define spin-aligned quasi-circular waveforms with parameters $\theta$ = $\{\mathcal{M}_{\mathrm{d}}$, $q$, $d_{\mathrm{L}}$, $\mathrm{dec}$, $\mathrm{ra}$, $\theta_{\mathrm{jn}}$, $\psi$, $t_{\mathrm{geocent}}$, $\chi_{1}$, $\chi_{2}$, $\phi_{c}\}$, where $\mathcal{M}_{\mathrm{d}}$, $q$, and $d_{\mathrm{L}}$ denote the detector-frame chirp mass, mass ratio, and luminosity distance, respectively; ra and dec denote the right ascension and declination of the source’s sky position, $\theta_{\mathrm{jn}}$ is the angle between the the line of sight and the total angular momentum $\vec{J}$, $\psi$ is the polarization angle, $t_{\mathrm{geocent}}$ is the coalescence time at the center of the Earth, and $\phi_{c}$ is the phase at coalescence. The waveforms also depend on the aligned spin parameters $\chi_{1}$ and $\chi_{2}$, defined as $\chi_{i}=a_{i}\cos\theta_{i}$, where $a_{i}=|\chi_{i}|$ is the spin magnitude and $\theta_{i}$ is the angle with respect to the orbital angular momentum, defined as:
[…]
We use frequency-domain waveforms, fixing the reference frequency to $f_{\mathrm{ref}}=20$ Hz (Varma:2021csh), with the waveform approximant IMRPhenomXPHM (PhysRevD.103.104056), which models subdominant modes $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4)]$ in addition to the dominant mode $(\ell,|m|)=(2,2)]$. A detailed discussion on higher-order modes can be found in Sec. IV.3.